### v2

# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

Abstract
Equipped with the continuous representation capability of Multi-Layer Perceptron (MLP), Implicit Neural Representation (INR) has been successfully employed for Arbitrary-scale Super-Resolution (ASR). However, the limited receptive field of the linear layers in MLP restricts the representation capability of INR, while it is computationally expensive to query the MLP numerous times to render each pixel. Recently, Gaussian Splatting (GS) has shown its advantages over INR in both visual quality and rendering speed in 3D tasks, which motivates us to explore whether GS can be employed for the ASR task. However, directly applying GS to ASR is exceptionally challenging because the original GS is an optimization-based method through overfitting each single scene, while in ASR we aim to learn a single model that can generalize to different images and scaling factors. We overcome these challenges by developing two novel techniques. Firstly, to generalize GS for ASR, we elaborately design an architecture to predict the corresponding image-conditioned Gaussians of the input low-resolution image in a feed-forward manner. Secondly, we implement an efficient differentiable 2D GPU/CUDA-based scale-aware rasterization to render super-resolved images by sampling discrete RGB values from the predicted contiguous Gaussians. Via end-to-end training, our optimized network, namely GSASR, can perform ASR for any image and unseen scaling factors. Extensive experiments validate the effectiveness of our proposed method. The project page can be found at https://mt-cly.github.io/GSASR.github.io/.
Figure 1: Comparison between GSASR and representative ASR methods on DIV2K. GSASR outperforms previous methods in visual quality with competitive inference time across both in-distribution ($\times$2, $\times$4) and out-of-distribution ($\times$6, $\times$12) scaling factors.
1 Introduction
Image Super-Resolution (SR) has been a challenging inverse problem in computer vision for many years, and a variety of deep learning-based (SR) methods have been developed in the past decade [61, 73, 39, 5, 57, 62, 71, 65, 66, 70, 6]. However, most of the existing SR models [79, 80, 60, 38, 77, 9, 16, 37, 74] are designed to operate only under integer or fixed scaling factors, i.e., ×2,×3,×4$\times$ 2$\times$ 3$\times$ 4× 2 , × 3 , × 4, due to the fact that they perform resolution upsampling by processing discrete pixels with de-convolutional layers [16] or pixel-shuffle layers [54]. As a result, it is hard for them to perform SR with decimal or even arbitrary scaling factors. Although the existing SR models could achieve arbitrary-scale zooming via adopting bicubic or bilinear interpolation, they would inevitably scarify image quality, leading to unsatisfied super-resolved results.
Inspired by the development of Implicit Neural Representation (INR) [48], Chen et al. presented LIIF [10] to address the problem of Arbitrary-scale SR (ASR) by making use of the continuity of Multi-Layer Perceptron (MLP) to query the RGB values of a high-resolution (HR) image and feeding its index to the MLP network. The following works try to improve the ASR performance with local feature estimators [32] or scale-aware attention modules [2]. Most of the existing INR-based ASR models [10, 32, 69, 2] rely heavily on MLP layers to predict RGB pixels. However, this implicit function lacks enough representation capacity, as the linear layers in MLP have a very limited receptive field, which leads to limited reconstruction performance. Moreover, since the computations of different indices are usually unshared during the MLP feed-forward process, existing INR-based ASR methods have low computational efficiency and difficulties in scaling up the model parameters.
Recently, Gaussian Splatting (GS) [26] has shown its advantages over INR-based methods, such as Neural Radiance Field (NeRF) [48], in terms of both accuracy and complexity, thanks to its high computational efficiency and powerful explicit representation capabilities. GS has been successfully applied to tasks such as 3D reconstruction [19, 4], representation [49, 18], editing [56, 11], generation [41, 12], etc. While some studies have explored the use of GS for low-level vision tasks, e.g., image deblurring [8, 81], image compression [78], video super-resolution [68], these models overfit on the repetitive images from multi-frames in a video clip or multi-view cameras, or overfit on a single image to complete the image compression task. None of these methods could be applied to single image SR tasks, since overfitting a model on a single image cannot meet the generalization requirements of SR, where a single model should be used to super-resolve any input low-resolution (LR) image to its high-resolution (HR) counterpart.
Instead of overfitting the model to a fixed scenario, we switch from the optimization-based manner of original 3D GS [26] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution), which adapts GS to the ASR task. Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [40, 79]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embedding to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [42] for them so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the contiguous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as condition. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
The advantages of our proposed GSASR are three-folds. (1) Due to the large receptive field brought by window-based self-attention modules, the Gaussian embeddings could sense long-range dependency from deep features and other embeddings. (2) The position of Gaussians are predicted explicitly, enabling Gaussians to concentrate on areas with complex textures, instead of being uniformly distributed over the whole image, thus leading to better reconstruction results. (3) Benefited from the developed efficient scale-aware 2D rasterization technique, GSASR is much faster than the state-of-the-art INR-based models. For example, under the scaling factor ×$\times 12$, GSASR could super-resolve an image to $720\times 720$ resolution within only $91$ ms, while the state-of-the-art INR-based model such as CiaoSR [2] will cost nearly $540$ ms.
Figure 2: Overview of GSASR. In the training phase, an LR image is fed into the encoder to extract image features, conditioned on which the learnable Gaussian embeddings are passed through the Condition Injection Block and Gaussian Interaction Block to output 2D Gaussians. These 2D Gaussians are then rendered into an SR image of a specified resolution through differential rasterization.
2 Related Work
SR aims to upscale the resolution of an LR input image. The community has witnessed a rapid development of SR methods, ranging from powerful feature extraction backbones [33, 40, 79, 80, 37, 21, 38, 47, 77, 9, 13, 35, 17, 74] to the utilization of generative models [31, 53, 60, 59, 76, 43, 44, 61, 73, 39, 34, 5, 51, 71, 66, 65, 57, 70, 62] for better perceptual quality. While different methods differ from each other on specific designs, they all rely on upsampling operators to magnify the resolution of the original LR image. Various upsampling strategies have been developed, ranging from hand-crafted interpolation operators, to deep neural network-based convolutional layers [16, 54], and to implicit neural representations [10].
Fixed-scale SR Methods. In the early stage, deep neural network-based SR models [15, 27, 28, 30] usually adopt interpolation operators to magnify the resolution. They first upsample the resolution of an LR image to the size of its HR counterpart, and then feed it into network for detail enhancement. Such a scheme is straightforward but suffers from high computational cost. FSRCNN [16] is proposed to apply de-convolutional layers to upsample the resolution of the deep features at the end of the network, so as to accelerate the inference speed. ESPCN [54] performs upsampling by rearranging the LR feature maps into a higher-resolution space through pixel-shuffle layers. However, since the pixels are discrete in feature domain, the de-convolution or pixel-shuffle layers cannot perform continuous sampling to reconstruct SR images with arbitrary scales. Due to these limitations, most of the existing SR methods are conducted on integer scaling factors.
Arbitrary-scale SR Methods. The interpolation operators can be used to continuously predict neighborhood pixels at any magnification factors, yet they will sacrifice the image quality and obtain blurry results. Meta-SR [24] equips the network with the arbitrary-scale magnification ability by designing a meta-upscaling module that leverages DNN to predict HR details from LR neighborhood embeddings. Wang et al. [58] performed arbitrary upsampling by introducing dynamic filters generated based on the scaling factor. Inspired by the success of INR-based methods [48] in 3D tasks, LIIF [10] creatively introduces INR to complete the ASR task by learning an implicit mapping function between a pixel’s RGB value and its index. Following LIIF, many INR-based models have been proposed. ITSRN [67] makes use of a transformer block and implicit position embedding to predict the super-resolved images. LTE [32] employs an implicit function to describe image features in Fourier domain. Based on Neural Operator [20], SRNO [64] adopts a kernel integral operator to fit the implicit function. BTC [50] employs B-spline implicit representation to achieve arbitrary-scale SR of screen content images. LINF [69] estimates a local implicit function with normalizing flow to model the naturalness of textures. CLIT [7] employs a cascaded cross-scale attention block with frequency encoding layers to sense neighborhood features. CiaoSR [2] introduces scale-aware attention-in-attention modules to extend the receptive field to promote the ASR performance. LMF [22] introduces a latent modulated function to compress the feature and accelerate the rendering process at the price of representation accuracy.
Very recently, GaussianSR [23] was proposed to employ GS for the ASR task. However, it has some significant drawbacks. Firstly, it only utilizes the position of RGB pixels from the low-resolution image to substite the center positions of Gaussians, heavily limiting the representation capability of Gaussians and leading to poor reconstruction performance. Secondly, it lacks a differentiable GPU/CUDA-based operator to complete the rasterization process, leading to low-efficiency and high computational cost. In contrast, our proposed GSASR predicts all of the parameters of Gaussians in an end-to-end manner, and we present an efficient differentiable 2D GPU/CUDA-based scare-aware rasterization to largely reduce the computational cost, leading to an effective and efficient model for accomplishing the ASR task.
3 Method
In this section, we first describe the concept and formulation of 2D Gaussian Spatting (GS), then present the model architecture and training details of our GSASR method.
3.1 2D GS Representation and Rasterization
2D GS Representation. Our key insight is to represent an image using contiguous Gaussians, which can fit various frequency components of an image due to the powerful expressive properties of GS. Specifically, considering an LR image $I_{LR}$ of size $H\times W$, we represent it using $N$ 2D Gaussians, denoted by $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$. Each Gaussian maintains properties of opacity $\alpha~{}\in[0,1]^{1}$, center position $\mu=\{\mu_{x}~{}\in\mathbb{R}^{1},\mu_{y}~{}\in\mathbb{R}^{1}\}$, standard deviation $\sigma=\{\sigma_{x}~{}\in(0,\infty)^{1},\sigma_{y}~{}\in(0,\infty)^{1}\}$, correlation coefficient $\rho~{}\in[-1,1]^{1}$, and normalized peak RGB value $c~{}\in[0,1]^{3}$. We define the color at position $(x,y)$ contributed by Gaussian $G$ as $G(x,y)$ (here we omit the index of Gaussians for simplicity of expression):
$G(x,y)=\alpha\cdot c\cdot f(x,y).$ (1)
The calculation of $f(x,y)$ is defined following typical 2D Gaussian distribution:
$f(x,y)=e^{-\frac{1}{2(1-\rho^{2})}(\frac{\Delta x^{2}}{\sigma_{x}^{2}}-\frac{2\rho\Delta x\Delta y}{\sigma_{x}\sigma_{y}}+\frac{\Delta y^{2}}{\sigma_{y}^{2}})},$ (2)
where $\Delta x=x-\mu_{x}$ and $\Delta y=y-\mu_{y}$. The color of the LR image at position $(x\in[0,W-1],y\in[0,H-1])$ is represented as the summarization of all $N$ Gaussians:
$I_{LR}(x,y)=\sum_{i=1}^{N}G_{i}(x,y).$ (3)
Gaussian functions are contiguous, enabling sampling at arbitrary positions. With a given scaling factor $s$, to obtain an SR image $I_{SR}$ of size $sH\times sW$, we sample the color at position $(x\in[0,sW-1],y\in[0,sH-1])$ with the following scale-aware rasterization:
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}).$ (4)
Fast Rasterization. Simply rendering an SR image by querying each pixel from all 2D Gaussians leads to a complexity of $\mathcal{O}(s^{2}HWN)$, which is too high for high-resolution images. Actually, a Gaussian generally focuses on a limited area and its contribution to pixel values decays rapidly with the increase of distance. Therefore, we introduce a rasterization ratio $r<1$ to control the rendering range of each Gaussian. Specifically, we handle all Gaussians in parallel and only render the pixels that are close enough to the Gaussian centers, greatly reducing the computational complexity to $\mathcal{O}(r^{2}s^{2}HWN)$ and making our algorithm practical. ($r$ is set to 0.1 in our implementation.) Note that our rasterization process is differential, which can be seamlessly integrated with neural networks for end-to-end optimization. Algorithm 1 describes the detailed steps of our scale-aware rasterization process. This algorithm is implemented via CUDA C++, which is GPU-friendly and achieves faster speed and low memory requirements.
1
2
Input: $N$ 2D Gaussians $\{G_{1},G_{2},\ldots,G_{N}\}$; LR image of size ($H$, $W$); scale factor $s$; rasterization ratio $r$.
Output: Rendered image $I_{SR}$
3
4Initialize $I_{SR}$ as an ($sH$, $sW$, 3) array of zeros.
5For each $G_{i}$ in $\{G_{1},G_{2},\ldots,G_{N}\}$ do
6       Initialize $\alpha$, $\mu_{x}$, $\mu_{y}$, $\sigma_{x}$, $\sigma_{y}$, $\rho$, $c$ from $G_{i}$.
7       For each pixel $(x,y)$ in $I_{SR}$ do
8             If $|x-\mu_{x}|<rsH$ and $|y-\mu_{y}|<rsW$ then
9                   Obtain $f(x/s,y/s)$ using Eq. 2;
10                   Obtain $G_{i}(x/s,y/s)$ using Eq. 1;
11                   $I_{SR}(x,y;s)\mathrel{+}=G_{i}(x/s,y/s)$.
12             end
13
14       end
15
16 end
Algorithm 1 2D GS Scale-aware Rasterization
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [40] as the feature extraction encoder. We test on DIV2K and LSDIR [36] datasets [55] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. The PSNR and SSIM indices are computed in the Y channel of Ycbcr space.
Scale Metrics Backbone: EDSR-baseline
Testing Dataset: DIV2K Testing Dataset: LSDIR
Meta
-SR
LIIF LTE SRNO LINF LMF
Ciao
-SR
Gaussian
-SR
GSASR
Meta
-SR
LIIF LTE SRNO LINF LMF
Ciao
-SR
Gaussian
-SR
GSASR
$\times$2 PSNR 36.02 36.05 36.10 36.27 36.21 36.21 36.42 36.10 36.65 31.36 31.43 31.50 31.65 31.49 31.61 31.78 31.47 32.14
SSIM 0.9455 0.9458 0.9461 0.9474 0.9461 0.9469 0.9476 0.9459 0.9495 0.9162 0.9170 0.9177 0.9197 0.9175 0.9193 0.9208 0.9174 0.9251
LPIPS 0.0889 0.0879 0.0869 0.0833 0.0887 0.0832 0.0835 0.0888 0.0767 0.0984 0.0963 0.0947 0.0916 0.0973 0.0916 0.0890 0.097 0.0823
DISTS 0.0571 0.0567 0.0567 0.0547 0.0564 0.0555 0.0543 0.0569 0.0514 0.0692 0.0688 0.0682 0.0662 0.0686 0.0667 0.0649 0.0685 0.0612
$\times$3 PSNR 32.34 32.38 32.43 32.51 32.48 32.52 32.61 32.42 32.89 27.88 27.93 27.99 28.09 27.98 28.07 28.12 27.97 28.46
SSIM 0.8877 0.8887 0.8892 0.8902 0.8892 0.8910 0.8911 0.8888 0.8960 0.8292 0.8312 0.8321 0.8345 0.8317 0.8347 0.8353 0.8314 0.8443
LPIPS 0.1977 0.1935 0.1915 0.1903 0.1945 0.1859 0.1873 0.1940 0.1782 0.2200 0.2128 0.2110 0.2075 0.2148 0.2056 0.2043 0.2142 0.1929
DISTS 0.1036 0.1035 0.1032 0.1015 0.1031 0.1414 0.0984 0.1041 0.0963 0.1255 0.1250 0.1242 0.1219 0.1246 0.1219 0.1174 0.1256 0.1155
$\times$4 PSNR 33.36 30.43 30.47 30.57 30.51 30.56 30.67 30.46 30.89 26.13 26.21 26.26 26.36 26.25 26.33 26.42 26.23 26.65
SSIM 0.8367 0.8388 0.8395 0.8415 0.8396 0.8416 0.8431 0.8389 0.8486 0.7577 0.7614 0.7627 0.7666 0.7621 0.7656 0.7681 0.7615 0.7774
LPIPS 0.2723 0.2662 0.2647 0.2616 0.2680 0.2607 0.2585 0.2684 0.2518 0.3074 0.2978 0.2957 0.2899 0.2998 0.2921 0.2865 0.3007 0.2777
DISTS 0.1394 0.1403 0.1397 0.1384 0.1401 0.1379 0.1370 0.1406 0.1301 0.1666 0.1678 0.1664 0.1647 0.1675 0.1647 0.1631 0.1679 0.1554
$\times$6 PSNR 28.07 28.20 28.26 28.34 28.26 28.31 28.38 28.19 28.60 24.29 24.38 24.43 24.50 24.41 24.47 24.54 24.38 24.71
SSIM 0.7618 0.7669 0.7679 0.7704 0.7674 0.7699 0.7711 0.7648 0.7784 0.6612 0.6684 0.6696 0.6736 0.6685 0.6722 0.6753 0.6654 0.6852
LPIPS 0.3696 0.3594 0.3657 0.3610 0.3646 0.3606 0.3547 0.3728 0.3435 0.4212 0.4080 0.4162 0.4088 0.4124 0.4115 0.3978 0.4226 0.3879
DISTS 0.1938 0.1954 0.1954 0.1939 0.1952 0.1934 0.1934 0.1969 0.1835 0.2247 0.2265 0.2256 0.2229 0.2262 0.2236 0.2217 0.2274 0.2120
$\times$8 PSNR 26.72 26.87 26.93 27.00 26.91 26.97 27.04 26.76 27.22 23.23 23.32 23.37 23.43 23.34 23.40 23.47 23.27 23.58
SSIM 0.7135 0.7207 0.7218 0.7243 0.7207 0.7235 0.7256 0.7155 0.7321 0.6032 0.6123 0.6137 0.6171 0.6118 0.6156 0.6194 0.6059 0.6269
LPIPS 0.4365 0.4212 0.4321 0.4261 0.4285 0.4264 0.4169 0.4445 0.4077 0.5020 0.4812 0.4939 0.4848 0.4868 0.4883 0.4703 0.5068 0.4611
DISTS 0.2305 0.2337 0.2348 0.2330 0.2341 0.2328 0.2314 0.2386 0.2214 0.2643 0.2673 0.2671 0.2640 0.2677 0.2654 0.2619 0.2713 0.2518
$\times$12 PSNR 25.06 25.21 25.28 25.34 25.23 25.31 25.37 24.96 25.50 22.00 22.08 22.13 22.17 22.09 22.15 22.20 21.95 22.26
SSIM 0.6595 0.6685 0.6693 0.6714 0.6678 0.6706 0.6731 0.6596 0.6777 0.5421 0.5525 0.5533 0.5558 0.5511 0.5542 0.5584 0.5424 0.5632
LPIPS 0.5290 0.5145 0.5300 0.5233 0.5249 0.5246 0.5080 0.5619 0.4975 0.6018 0.5882 0.6047 0.5953 0.5962 0.6013 0.5750 0.6363 0.5642
DISTS 0.2852 0.2905 0.2938 0.2914 0.2934 0.2916 0.2885 0.3038 0.2787 0.3210 0.3256 0.3267 0.3238 0.3280 0.3255 0.3203 0.3366 0.3106
$\times$16 PSNR 24.00 24.13 24.20 24.25 24.14 24.23 24.27 23.80 24.38 21.22 21.29 21.34 21.38 21.31 21.32 21.40 21.12 21.42
SSIM 0.6307 0.6402 0.6407 0.6423 0.6394 0.6418 0.6443 0.6306 0.6473 0.5120 0.5221 0.5227 0.5245 0.5209 0.5213 0.5271 0.5122 0.5296
LPIPS 0.5924 0.5754 0.5912 0.5847 0.5876 0.5861 0.5666 0.6325 0.5563 0.6687 0.6529 0.6691 0.6617 0.6622 0.6738 0.6386 0.7101 0.6280
DISTS 0.3297 0.3357 0.3398 0.3367 0.3412 0.3371 0.3336 0.3582 0.3242 0.3638 0.3683 0.3701 0.3671 0.3727 0.3687 0.3636 0.3866 0.3538
$\times$18 PSNR 23.57 23.71 23.77 23.83 23.71 23.81 23.85 23.37 23.93 20.92 20.99 21.04 21.07 21.00 21.03 21.09 20.81 21.11
SSIM 0.6212 0.6306 0.6309 0.6323 0.6297 0.6319 0.6344 0.6213 0.6367 0.5027 0.5124 0.5126 0.5142 0.5112 0.5115 0.5165 0.503 0.5185
LPIPS 0.6055 0.5966 0.6130 0.6066 0.6083 0.6081 0.5881 0.6563 0.5790 0.6807 0.6749 0.6916 0.6837 0.6828 0.6948 0.6609 0.7343 0.6514
DISTS 0.3490 0.3540 0.3584 0.3555 0.3606 0.3557 0.3518 0.3813 0.3419 0.3812 0.3854 0.3875 0.3843 0.3905 0.3852 0.3807 0.4071 0.3710
$\times$24 PSNR 22.63 22.74 22.81 22.84 22.74 22.83 22.87 22.41 22.90 20.25 20.31 20.35 20.39 20.32 20.34 20.40 20.11 20.38
SSIM 0.6027 0.6108 0.6112 0.6119 0.6101 0.6117 0.6141 0.6037 0.6150 0.4854 0.4938 0.4938 0.4845 0.4930 0.4924 0.4967 0.4862 0.4973
LPIPS 0.6512 0.6444 0.6600 0.6539 0.6545 0.6561 0.6352 0.7042 0.6299 0.7234 0.7214 0.7374 0.7292 0.7261 0.7421 0.7067 0.7812 0.7028
DISTS 0.4022 0.4009 0.4063 0.4026 0.4106 0.4036 0.3980 0.4380 0.3877 0.4292 0.4274 0.4308 0.4269 0.4348 0.4297 0.4222 0.4579 0.4117
$\times$30 PSNR 21.97 22.07 22.12 22.16 22.07 22.03 22.18 21.75 22.19 19.75 19.80 19.84 19.86 19.81 19.83 19.87 19.60 19.82
SSIM 0.5982 0.5998 0.6000 0.6003 0.5994 0.5966 0.6021 0.5943 0.6025 0.4754 0.4824 0.4823 0.4824 0.4817 0.4818 0.4839 0.4766 0.4842
LPIPS 0.7018 0.6773 0.6914 0.6862 0.6859 0.7137 0.6682 0.7317 0.6648 0.7386 0.7537 0.7676 0.7610 0.7575 0.7645 0.7284 0.8084 0.7368
DISTS 0.4474 0.4375 0.4445 0.4400 0.4500 0.4476 0.4339 0.4821 0.4232 0.4702 0.4603 0.4645 0.4597 0.4702 0.4608 0.4535 0.4986 0.4417
3.2 GSASR: Architecture and Modules
Conventional 3D Gaussian Splatting (GS) methods [26] directly optimize 3D Gaussians to fit a specific scene. As a result, they lack the generalization capability for the ASR tasks, where the model should be generalized to unseen images. To overcome this limitation, we introduce a novel network architecture, which conditions 2D Gaussians on low-resolution (LR) input images, generating content-aware Gaussian representations with enhanced generalization capabilities. Our proposed method, namely GSASR, is illustrated in Figure 2. It begins with an off-the-shelf encoder that extracts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ from the LR input $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, serving as input to the subsequent Gaussian generation process. The key of GSASR lies in the proposed Gaussian decoder, which receives image feature $\mathbf{F}$ and a learnable Gaussian embedding as inputs, and outputs 2D Gaussians $\mathcal{G}$ to depict the input image in contiguous Gaussian space. Our Gaussian decoder consists of a condition injection block, a Gaussian interaction block, and a Gaussian primary head. Finally, the output of Gaussian decoder is rasterized into the SR image.
Gaussian Embedding. In the GS decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. A straightforward idea is to maintain $N$ independent learnable Gaussian embeddings that interact with image features to output $\mathcal{G}$. However, this approach suffers from both effectiveness and computational efficiency. First, the LR images exhibit variable sizes in the test phase. Rendering a fixed number of 2D Gaussians is inadequate to represent the details of a high-resolution LR image. Second, maintaining $N$ independent Gaussian embeddings not only introduces a significant memory burden but also faces optimization challenges, primarily attributed to the initial homogeneity and independent optimization of embeddings.
To address these limitations, we propose two key innovations. First, instead of setting a fixed value for $N$, we let $N$ be proportional to the LR image size ($N\propto H\times W$), employing more Gaussians to represent larger inputs. All Gaussian embeddings share the same initial weights, as shown in Figure 2. Second, to mitigate the optimization complexity and computational demands, inspired by Deform-DETR [82], we assign the $N$ Gaussian embeddings to different reference positions $p\in\mathbb{R}^{N\times 2}$ during initialization, which is obtained by sampling $N$ points at equal intervals from an $H\times W$ image. In this way, different Gaussian embeddings interpret the content at different regions based on their assigned reference positions, while they share the same weight initialization.
Figure 3: Demonstration of the sampling process in 2D rasterization. A larger upsampling scale factor employs smaller sampling intervals from continuous Gaussians.
Condition Injection Block. As shown in Figure 2, we utilize cross-attention to incorporate feature $\mathbf{F}$ into Gaussian embeddings $\mathbf{E}$. Such a technique is widely used in models like Stable Diffusion [52]. To reduce the computational complexity, we perform window-based cross-attention [42] only in pre-defined local windows as follows:
$\text{Attention}(Q,K,V)=\text{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V,$ (5)
where $Q$ is from Gaussian embedding $\mathbf{E}$, $K$ and $V$ are from image feature $\mathbf{F}$. The $i$-th embedding $\mathbf{E}_{i}$ at position $p_{i}$ interacts with the $j$-th pixel features within the window. $B$ is the relative position bias in the feature map.
Table 2: Quantitative comparison between representative ASR models and our GSASR. All models use the same RDN [80] backbone as the feature extraction encoder. We test on DIV2K and LSDIR [36] datasets [55] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. The PSNR and SSIM indices are computed in the Y channel of Ycbcr space.
Scale Metrics Backbone: RDN
Testing Dataset: DIV2K Testing Dataset: LSDIR
Meta
-SR
LIIF LTE SRNO LINF LMF
Ciao
-SR
Gaussian
-SR
GSASR
Meta
-SR
LIIF LTE SRNO LINF LMF
Ciao
-SR
Gaussian
-SR
GSASR
$\times$2 PSNR 36.54 36.37 36.41 36.53 36.53 36.48 36.67 36.54 36.73 31.97 31.87 31.96 32.10 31.94 32.02 32.16 31.95 32.26
SSIM 0.9483 0.9481 0.9484 0.9493 0.9483 0.9487 0.9494 0.9484 0.9500 0.9224 0.9221 0.9228 0.9247 0.9223 0.9234 0.9247 0.9228 0.9261
LPIPS 0.0816 0.0828 0.0816 0.0783 0.0821 0.0806 0.0797 0.0829 0.0757 0.0872 0.0876 0.0861 0.0832 0.0874 0.0866 0.0839 0.0875 0.0812
DISTS 0.0542 0.0543 0.0535 0.0528 0.0541 0.0540 0.0524 0.0543 0.0509 0.0644 0.0645 0.0635 0.0623 0.0645 0.0639 0.0615 0.0640 0.0603
$\times$3 PSNR 32.79 32.68 32.73 32.83 32.77 32.78 32.89 32.74 32.97 28.34 28.26 28.34 28.45 28.30 28.37 28.46 28.33 28.56
SSIM 0.8935 0.8934 0.8938 0.8951 0.8936 0.8947 0.8952 0.8939 0.8970 0.8399 0.8398 0.8409 0.8435 0.8398 0.8421 0.8432 0.8407 0.8462
LPIPS 0.1850 0.1842 0.1837 0.1817 0.1846 0.1807 0.1821 0.1847 0.1768 0.1997 0.1983 0.1979 0.1947 0.1995 0.1954 0.1953 0.1993 0.1908
DISTS 0.1000 0.0998 0.0987 0.0980 0.0999 0.0984 0.0965 0.0998 0.0955 0.1192 0.1188 0.1176 0.1159 0.1194 0.1169 0.1139 0.1189 0.1139
$\times$4 PSNR 30.78 30.71 30.75 30.85 30.77 30.81 30.91 30.76 30.96 26.53 26.48 26.54 26.64 26.51 26.58 26.66 26.53 26.73
SSIM 0.8455 0.8449 0.8459 0.8478 0.8453 0.8467 0.8481 0.8457 0.8500 0.7724 0.7714 0.7734 0.7767 0.7719 0.7744 0.7770 0.7727 0.7801
LPIPS 0.2568 0.2566 0.2558 0.2525 0.2574 0.2546 0.2525 0.2570 0.2505 0.2831 0.2838 0.2822 0.2771 0.2840 0.2817 0.2768 0.2837 0.2752
DISTS 0.1341 0.1354 0.1341 0.1330 0.1355 0.1341 0.1327 0.1347 0.1288 0.1581 0.1603 0.1589 0.1567 0.1609 0.1584 0.1563 0.1595 0.1533
$\times$6 PSNR 28.40 28.44 28.49 28.57 28.48 28.52 28.61 28.46 28.65 24.55 24.59 24.64 24.70 24.61 24.66 24.73 24.60 24.77
SSIM 0.7715 0.7739 0.7749 0.7772 0.7739 0.7757 0.7777 0.7727 0.7800 0.6756 0.6791 0.6806 0.6841 0.6787 0.6815 0.6853 0.6772 0.6881
LPIPS 0.3505 0.3468 0.3551 0.3510 0.3523 0.3534 0.3452 0.3584 0.3443 0.3960 0.3915 0.4020 0.3964 0.3961 0.4009 0.3864 0.4043 0.3867
DISTS 0.1837 0.1882 0.1878 0.1861 0.1890 0.1873 0.1867 0.1880 0.1817 0.2110 0.2166 0.2161 0.2131 0.2181 0.2154 0.2131 0.2167 0.2093
$\times$8 PSNR 26.97 27.07 27.13 27.20 27.11 27.16 27.24 26.99 27.28 23.44 23.49 23.54 23.60 23.52 23.57 23.63 23.45 23.63
SSIM 0.7222 0.7274 0.7286 0.7308 0.7272 0.7293 0.7317 0.7228 0.7339 0.6151 0.6221 0.6230 0.6264 0.6211 0.6240 0.6281 0.6156 0.6299
LPIPS 0.4133 0.4072 0.4214 0.4166 0.4163 0.4188 0.4075 0.4290 0.4060 0.4679 0.4627 0.4798 0.4729 0.4711 0.4772 0.4592 0.4875 0.4593
DISTS 0.2181 0.2251 0.2268 0.2248 0.2273 0.2260 0.2244 0.2285 0.2192 0.2489 0.2562 0.2569 0.2537 0.2588 0.2563 0.2528 0.2601 0.2490
$\times$12 PSNR 25.23 25.39 25.45 25.52 25.41 25.48 25.55 25.12 25.56 22.13 22.21 22.26 22.31 22.22 22.28 22.33 22.06 22.30
SSIM 0.6653 0.6739 0.6748 0.6768 0.6730 0.6754 0.6783 0.6648 0.6794 0.5494 0.5597 0.5605 0.5629 0.5576 0.5608 0.5651 0.5484 0.5656
LPIPS 0.5001 0.4981 0.5180 0.5132 0.5120 0.5153 0.4982 0.5455 0.4947 0.5671 0.5685 0.5899 0.5823 0.5803 0.5884 0.5635 0.6174 0.5621
DISTS 0.2703 0.2809 0.2857 0.2835 0.2855 0.2844 0.2811 0.2921 0.2759 0.3055 0.3138 0.3165 0.3133 0.3187 0.3160 0.3116 0.3252 0.3077
$\times$16 PSNR 24.12 24.30 24.36 24.42 24.30 24.38 24.44 23.93 24.44 19.34 21.40 21.45 21.49 21.41 21.45 21.50 21.20 21.47
SSIM 0.6346 0.6446 0.6452 0.6467 0.6433 0.6456 0.6484 0.6340 0.6487 0.4813 0.5273 0.5277 0.5294 0.5253 0.5269 0.5318 0.5156 0.5318
LPIPS 0.5606 0.5580 0.5788 0.5739 0.5751 0.5760 0.5572 0.6166 0.5537 0.6762 0.6343 0.6556 0.6487 0.6488 0.6598 0.6281 0.6932 0.6261
DISTS 0.3117 0.3243 0.3312 0.3288 0.3324 0.3295 0.3268 0.3455 0.3211 0.3845 0.3566 0.3605 0.3575 0.3638 0.3596 0.3562 0.3745 0.3512
$\times$18 PSNR 23.69 23.87 23.93 23.99 23.87 23.96 24.01 23.48 23.98 20.99 21.09 21.14 21.17 21.10 21.13 21.18 20.88 21.14
SSIM 0.6244 0.6346 0.6349 0.6362 0.6333 0.6353 0.6380 0.6239 0.6379 0.5056 0.5165 0.5168 0.5181 0.5147 0.5160 0.5205 0.5056 0.5203
LPIPS 0.5776 0.5805 0.6013 0.5968 0.5957 0.5987 0.5796 0.6418 0.5766 0.6489 0.6575 0.6490 0.6720 0.6696 0.6822 0.6511 0.7188 0.6497
DISTS 0.3303 0.3419 0.3499 0.3476 0.3513 0.3480 0.3455 0.3676 0.3391 0.3633 0.3733 0.3784 0.3752 0.3814 0.3764 0.3740 0.3944 0.3682
$\times$24 PSNR 22.71 22.87 22.92 22.97 22.87 22.93 22.99 22.49 22.96 19.26 20.39 20.44 20.47 20.41 20.43 20.48 20.16 20.42
SSIM 0.6041 0.6139 0.6138 0.6145 0.6128 0.6136 0.6166 0.6048 0.6162 0.4728 0.4967 0.4965 0.4973 0.4953 0.4955 0.4995 0.4873 0.4985
LPIPS 0.6273 0.6301 0.6505 0.6461 0.6423 0.6487 0.6292 0.6939 0.6275 0.7112 0.7073 0.7274 0.7207 0.7143 0.7316 0.7005 0.7701 0.7006
DISTS 0.3830 0.3865 0.3978 0.3955 0.4008 0.3961 0.3932 0.4224 0.3851 0.4205 0.4145 0.4220 0.4185 0.4260 0.4212 0.4172 0.4434 0.4095
$\times$30 PSNR 17.25 22.18 22.23 22.27 32.18 22.18 22.28 21.81 22.23 19.78 19.86 19.91 19.93 19.89 19.90 19.94 19.64 19.86
SSIM 0.5546 0.6016 0.6017 0.6024 0.6010 0.5995 0.6041 0.5947 0.6032 0.4758 0.4842 0.4840 0.4844 0.4835 0.4837 0.4861 0.4770 0.4851
LPIPS 0.6817 0.6652 0.6833 0.6799 0.6757 0.7009 0.6627 0.7247 0.6633 0.7164 0.7418 0.7598 0.7540 0.7468 0.7570 0.7269 0.8009 0.7358
DISTS 0.4405 0.4215 0.4359 0.4332 0.4394 0.4361 0.4301 0.4650 0.4213 0.4509 0.4454 0.4562 0.4520 0.4605 0.4536 0.4495 0.4823 0.4407
Gaussian Interaction Block. In the condition injection block, each Gaussian embedding $\mathbf{E}_{i}$ is processed independently, resulting in a lack of information exchange among different Gaussian embeddings. To address this issue, we stack $L$ Gaussian interaction blocks to enhance the interaction among Gaussian embeddings. The Gaussian interaction block is designed based on Swin-transformer architecture, which iteratively stacks the window attention layers and shifted window attention layers. In attention layers, we perform self-attention for embeddings whose reference positions are located in the same window, similar to the condition injection block. Besides, we introduce a cross-attention layer to make embeddings scale-aware so that the Gaussian embeddings can learn to adjust the properties according to the given scale factor.
Meta-SR
LMF
LIIF
CiaoSR
LTE
GaussianSR
SRNO
GSASR (Ours)
LINF
GT
Figure 4: Visualization of GSASR and the competing methods under ×$\times 4$ scaling factor with EDSR [40] feature extraction backbone. The competing methods result in blurry details, while GSASR generates much clear textures.
Gaussian Primary Head. We employ several disentangled property heads to convert Gaussian embeddings $\mathbf{E}$ into 2D Gaussians $\mathcal{G}$. The $i$-th Gaussian embedding $\mathbf{E}_{i}$ is fed into a lightweight MLP to output its opacity $\alpha_{i}$, color $c_{i}$, stand deviation $\sigma_{i}$, reference offset $o_{i}$, and correlation $\rho_{i}$. Note that the position is obtained by adding the reference position and predicted offset $\mu_{i}=p_{i}+o_{i}$. The five property heads are decoupled so that different heads can be adapted to predict specific properties. We apply the tanh function to ensure $\rho\in[-1,1]$, and use the sigmoid function for $\{\alpha,c,\sigma\}$ to ensure that their values are physically meaningful. No activation function is applied to $o$ since the positions of Gaussians in 2D images are expected to be free.
Meta-SR
LMF
LIIF
CiaoSR
LTE
GaussianSR
SRNO
GSASR (Ours)
LINF
GT
Figure 5: Visualization of GSASR and the competing methods under ×$\times 6$ scaling factor with RDN [80] feature extraction backbone. The competing methods produce blurry textures, while GSASR generates much clearer contents.
Rendering. In Eq. 4, the rasterization determines the sampling interval of the 2D Gaussians based on the scale factor. As illustrated in Figure 3, larger upsampling factors use smaller sampling intervals, while the smaller upsampling factors use larger intervals. Due to the excellent representational capacity of Gaussians, high-quality sampling results can be achieved even in areas with complex textures. After rendering the SR image from rasterization, it is compared with the HR ground truth to calculate the loss $L_{1}$.
4 Experiments
4.1 Experimental Setup
Training Details. We utilize the widely-used DIV2K [55] dataset as training set. We follow [10, 2] to fix LR patches to $48\times 48$. The scaling factor $s$ is randomly selected from $[1.0,4.0]$. We firstly crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply the bicubic down-sampling [2] to obtain the corresponding LR patches. We employ EDSR [40] and RDN [80] as the image encoder backbones to extract feature from the input LR images. We train our GSARS on $4$ NVIDIA A100 GPUs for $500,000$ iterations, with a batch size of 64. The initial learning rate is set to $2e^{-4}$, and halves at $250,000$, $400,000$, $450,000$, $475,000$ iterations. The Adam [29] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to $2,000$.
Table 3: Comparison of computational costs. We report the results using EDSR-backbone [40] as image encoder. Apart from the PSNR/SSIM/LPIPS/DISTS metrics, we report the average inference time (ms) and the GPU memory usage (MB). The best results are highlighted in red.
Scale Computational Cost and Performance Backbone: EDSR
Testing GT Size: 720 * 720
Meta-SR LIIF LTE SRNO LINF LMF CiaoSR GaussianSR GSASR
$\times$2 PSNR 37.09 37.14 37.16 37.34 37.42 37.29 37.59 37.18 37.84
SSIM 0.9475 0.9487 0.9491 0.9501 0.9490 0.9498 0.9502 0.9489 0.9522
LPIPS 0.0787 0.0777 0.0768 0.0743 0.0785 0.0736 0.0744 0.0785 0.0676
DISTS 0.0741 0.0737 0.0736 0.0713 0.0731 0.0724 0.0710 0.0738 0.0676
Inference Time 186 454 126 107 86 209 2372 754 1573
GPU Memory 670 548 490 6301 3573 4005 49152 5200 13367
$\times$3 PSNR 33.39 33.44 33.45 33.59 33.63 33.56 33.76 33.47 34.04
SSIM 0.8921 0.8931 0.8935 0.8953 0.8936 0.8953 0.8955 0.8932 0.9004
LPIPS 0.1813 0.1778 0.1756 0.1729 0.1783 0.1701 0.1718 0.1780 0.1625
DISTS 0.1258 0.1257 0.1255 0.1230 0.1255 0.1236 0.1201 0.1262 0.1182
Inference Time 161 438 118 114 89 147 2016 717 806
GPU Memory 493 570 333 6282 3412 1798 10002 5138 6000
$\times$4 PSNR 31.38 31.48 31.48 31.65 31.62 31.60 31.79 31.49 32.01
SSIM 0.8417 0.8437 0.8444 0.8472 0.8444 0.8466 0.8482 0.8437 0.8536
LPIPS 0.2545 0.2490 0.2472 0.2429 0.2506 0.2438 0.2416 0.2511 0.2344
DISTS 0.1638 0.1652 0.1646 0.1626 0.1653 0.1930 0.1617 0.1656 0.1542
Inference Time 47 182 114 95 66 68 1165 686 543
GPU Memory 432 308 279 6275 3357 1139 3331 5048 3420
$\times$6 PSNR 29.08 29.24 29.25 29.39 29.33 29.34 29.51 29.19 29.69
SSIM 0.7686 0.7735 0.7744 0.7777 0.7738 0.7766 0.7790 0.7715 0.7851
LPIPS 0.3455 0.3362 0.3420 0.3361 0.3411 0.3374 0.3298 0.3492 0.3223
DISTS 0.2179 0.2199 0.2195 0.2176 0.2201 0.2179 0.2171 0.2214 0.2083
Inference Time 46 176 118 94 65 55 731 692 265
GPU Memory 387 264 239 6270 3316 912 1548 5224 1578
$\times$8 PSNR 27.74 27.89 27.94 28.04 27.97 28.00 28.14 27.77 28.25
SSIM 0.7216 0.7284 0.7297 0.7326 0.7285 0.7314 0.7345 0.7235 0.7397
LPIPS 0.4132 0.3956 0.4062 0.3988 0.4025 0.4017 0.3892 0.4187 0.3810
DISTS 0.2543 0.2569 0.2573 0.2555 0.2576 0.2570 0.2542 0.2623 0.2448
Inference Time 42 170 112 96 64 64 621 666 195
GPU Memory 371 248 224 6269 3302 832 1503 5012 1051
$\times$12 PSNR 26.03 26.18 26.23 26.32 26.24 26.28 26.39 25.91 26.50
SSIM 0.6687 0.6773 0.6783 0.6805 0.6768 0.6796 0.6825 0.6688 0.6864
LPIPS 0.4990 0.4861 0.5011 0.4941 0.4958 0.4959 0.4769 0.5314 0.4701
DISTS 0.3063 0.3114 0.3141 0.3113 0.3140 0.3127 0.3084 0.3241 0.3007
Inference Time 41 172 113 91 64 52 540 688 91
GPU Memory 360 237 214 6268 3292 775 1470 5214 472
SR Image
Position Distribution of Gaussians
Colorized Gaussians
Figure 6: Demonstration of the rich expressiveness of Gaussians for ASR. The position distribution is obtained by setting $\{\sigma,\rho,c\}$ to fixed values. In the middle image, one could observe that Gaussians are evenly distributed in regions with simple textures or flatten area (such as the sky), while their positions are adjusted in regions with complex textures to fit details. (Please zoom in for better observation.) In the right image, we randomly select parts of Gaussians and highlight their colors to red. One can see that 2D Gaussians can learn to fit the different object shapes ($e.g.,$ the edge of window).
Comparison Methods. We compare our proposed GSASR with Meta-SR [24], LIIF [10], LTE [32], SRNO [64], LINF [69], CiaoSR [2], LMF [22], and GaussianSR [23].
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [72], DIV2K100 [55], Urban100 [25], BSDS100 [45], Manga109 [46], General100 [16], and LSDIR [36] datasets under scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. To validate the generalization on higher scaling factors, we further report the results with ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$ scaling factors on DIV2K100 [55] and LSDIR [36] datasets. For performance measurement, we report PSNR, SSIM [63], LPIPS [75], and DISTS [14] indices. The PSNR and SSIM are computed in the Y channel of Ycbcr space. For the comparison of computational cost, we crop $100$ GT patches with $720\times 720$ resolution from DIV2K [55], and use bicubic interpolation to generate the corresponding LR images with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. We report the average inference time (ms) and GPU memory usage (MB) on a single NVIDIA A100 GPU.
4.2 Experiment Results
Quantitative Results. Table 1 and Table 2 show the numerical results of our proposed GSASR and other comparison methods. One could see that GSASR outperforms existing methods in terms of both fidelity metrics (PSNR and SSIM [63]) and perceptual quality metrics (LPIPS [75] and DISTS [14]) under both the EDSR [40] and RDN [80] backbones and scaling factors from ×$\times 2$ to ×$\times 30$. It obtains significant improvements over the methods based on INR, exhibiting greater potential for ASR. Compared with the recent GaussianSR method [23], which also employs GS for feature representation, GSASR shows clear advantages. GSASR predicts the position of Gaussians via an end-to-end manner, while GaussianSR [23] simply replaces the position parameters with the fixed position of RGB pixels of the LR image and loses representation capability. Besides, GaussianSR lacks GPU/CUDA-based rasterization to accelerate the inference speed and it fixes the number of Gaussians to $100$, no matter how large the scaling factor is. Therefore, the performance of GaussianSR drops a lot under high magnification factors. On the contrary, we present an efficient 2D GPU/CUDA-based rasterization to embed much more Gaussians to promote ASR performance. As a result, although GSASR is trained with a scaling factor from $1.0$ to $4.0$, it could not only obtain excellent performance for in-distribution scaling factors (i.e., scaling factor is less than $4$), but also show strong capability in dealing with out-of-distribution situations (i.e., the scaling factor is larger than $4$). More results on Set5 [1], Set14 [72], Urban100 [25], BSDS100 [45], Manga109 [46], General100 [16] could be found in the Appendix.
Qualitative Results. Figure 4 and Figure 5 show the qualitative comparison of competing methods. We can see that the compared methods mostly produce blurry details, for example, the holes in Figure 4, while GSASR generates much clearer edges and details, owing to the strong representation ability from 2D Gaussians. More visualization results can be found in the Appendix. To explain why GSASR works well for the ASR task, we visualize the positions together with the shape of Gaussians in Figure 6. One could see that the positions of Gaussians tend to cluster in areas with complex textures (such as the windows), while they are uniformly distributed in flat areas (such as the sky). This validates that our proposed strategy is highly suitable for restoring complex texture details. Meanwhile, from the blue box and green box in Figure 6, one could find that the Gaussians will adaptively adjust their orientations and scales based on the direction and shape of the textures.
Computational Costs. Table 3 shows that GSASR surpasses all the competing methods in most of the fidelity (PSNR, SSIM) and perceptual (LPIPS, DISTS) measures. In terms of speed, GSASR is nearly twice as fast as the state-of-the-art method CiaoSR [2] under ×3,×4$\times$ 3$\times$ 4× 3 , × 4 scaling factors, and speed advantage becomes more significant when handling with larger scaling factors. Thanks to the efficient 2D GPU/CUDA-based rasterization, GSASR could render an image with high speed without sacrificing the reconstruction fidelity. The computational cost with RDN backbone could be found in the Appendix.
Ablation Study. We conduct ablation studies on the number of Gaussians $N$, and the functionality of the reference position. Details are in the Appendix.
5 Conclusion and Limitation
We presented GSASR, a brandly-new 2D GS-based ASR model. To adapt GS to the ASR task, we firstly elaborately designed an architecture to covert an LR input to image-conditioned Gaussians, then implemented an efficient differentiable scale-aware 2D GPU/CUDA-based rasterization to render images with both high quality and fast speed. Through sampling values from Gaussians, we could render an output image with arbitrary magnification factors. Extensive experiments demonstrated that our GSASR model has a much more powerful representation capability, together with more friendly computational costs, than Implicit Neural Function based methods.
Limitations. GSASR still has some limitations. First, its performance depends on the number of Gaussians, especially under ultra-high scaling factors. While more Gaussians could bring better performance, it will sacrifice the inference speed. Second, to convert an image into Gaussians, a large number of parameters are used to fit the complex mapping function. How to reduce the number of parameters needs further investigation.